A ball is thrown in the air from the top of a building. Its height, in meters above ground, as a function of time, in seconds, is given by h(t)= -4.9t^2+ 24t + 8 What is the height of the building? What is the maximum height reached by the ball? How long does it take to reach maximum height? Also, find a rigorous algebraic solution for the problem.


1
Expert's answer
2021-05-03T07:02:09-0400

Solution

The height of the building is given when t=0

thus h(0)= -4.9 (0)2 24 (0) + 8 = 8

Therefore, the height of the building = 8

At maximum height the first derivative of the function is equal to zero

hence h'(x)=-9.8t +24=0;

Thus time taken to reach the maximum height, t= 24/9.8 = 2.449.

Hence the maximum height reached by the ball = h(2.449) = -4.9 (2.449)2 +24 (2.449) + 8 = 37.3878